OpenStudy (amonoconnor):

I think this should be a super easy question, because I'm just getting a conflicting answer to what my book has, and I can't find an error in my algebra. Would someone check my answer? The problem: Evaluate integral[(5-2t+3t^2)dt] from 1 to 4. Any and all help is greatly appreciated!

OpenStudy (anonymous):

What expression did you get for the integral?

OpenStudy (amonoconnor):

Here's what I did: \[\int\limits_{1}^{4}(5-2t+3t^2)dt\] \[f(t) = 5 - 2t + 3t^2\] \[F(t) = 5t - t^2 + t^3\] \[\int\limits_{1}^{4}(5-2t+3t^2)dt\] =\[F(b) - F(a)\]= \[(5(4) - (4^2) + (4^3))-(5(1)-(1^2)+(1^3))\] = \[(20-16+64)-(5-1+1)\] \[(-4 + 64) - (4+1)\] \[60 - 5 \] = 55

OpenStudy (amonoconnor):

I see it... damn it!! My bad... I See my mistake...

OpenStudy (amonoconnor):

sometimes it takes writing it out, and asking a question to make your brain stop and restart...

OpenStudy (anonymous):

Simple math error. Sometimes writing out makes it as plain as day. Good job.

OpenStudy (amistre64):

algebra is usually the bane of calculus ... and on that rare occasion, arithmetic :)